# Resolving Spurious Multifractality in Discrete Systems: A Finite-Size Scaling Protocol for MFDFA in the 2D Ising Model

### 04 Mar 2026 v1

@@ VI.1 Resolution of the ”Spurious Multifractality” Controversy
VI.2 MFDFA as a Configural Diagnostic for Disorder
The contrast provided by the RBIM results (Fig. 5) is paramount for the validation of the method. If MFDFA were simply insensitive to subtle correlation structures, it would have failed to detect the complexity of the disordered model. The fact that the spectrum remains broad ($\Delta\alpha\approx 0.23$) for the RBIM in the same range of $L$ where the pure model collapses proves that MFDFA acts as a high-precision discriminator.
Physically, the RBIM is characterized by ”Griffiths phases”—rare regions devoid of impurities that order locally at temperatures different from the bulk $T_{c}$. These rare regions introduce a spatial distribution of local correlation lengths $\xi(\mathbf{r})$, which manifests mathematically as a true spectrum of singularities $f(\alpha)$. Therefore, we establish the spectral width $\Delta\alpha$ as an effective order parameter for distinguishing between ”clean” criticality (monofractal) and ”dirty” criticality (multifractal) in experimental data where the underlying Hamiltonian might be unknown.
@@ VI.4 Generality Beyond the 2D Ising Model
    Systems with quenched disorder (random field Ising, spin glasses, diluted models): Genuine broad multifractal spectra should persist, reflecting the spatial heterogeneity of local critical temperatures. Our RBIM results (Fig. 5) demonstrate the method’s sensitivity to this regime.
@@ VII Conclusion
Our comparative analysis with the Random Bond Ising Model provides crucial validation. The RBIM exhibits a broad, robust multifractal spectrum ($\Delta\alpha\approx 0.23$) that persists under FSS, demonstrating that our protocol is sufficiently sensitive to detect genuine multifractality when physically present. This establishes MFDFA not merely as a numerical tool, but as a sensitive diagnostic capable of distinguishing “clean” monofractal criticality from “dirty” disorder-dominated complexity. The spectral width $\Delta\alpha$ emerges as an effective order parameter for this distinction.

### 30 Mar 2026 v2

@@ VI.1 Resolution of the ”Spurious Multifractality” Controversy
VI.2 FSS as a Discriminator: Transient Artifacts versus Genuine Disorder
A fundamental question arises: is the severe finite-size distortion observed in the pure Ising model unique to perfectly ordered systems, or does it also afflict models with heterogeneous correlations? To address this, we apply our MFDFA Finite-Size Scaling protocol to the Random Bond Ising Model (RBIM). The RBIM introduces quenched bond disorder, leading to the formation of Griffiths phases—rare regions that order locally and introduce a broad distribution of correlation lengths.
Historically, analyses of single-size RBIM lattices have yielded broad multifractal spectra, seemingly confirming the complex nature of the disorder. Figure 5 presents our FSS analysis of the RBIM for system sizes $L\in[64,512]$, visualized through the generalized Hurst exponent $H(q)$. For a monofractal system, $H(q)$ is constant. Strikingly, we find that the broad dispersion of $H(q)$ observed at small scales ($L=64$) is not robust. As the system size increases, the $H(q)$ curves systematically flatten, and the generalized width $\Delta H=H(q_{min})-H(q_{max})$ continuously decays toward the asymptotic monofractal limit.
This comparative FSS analysis constitutes a crucial methodological finding with two complementary aspects. First, it demonstrates that even in systems with genuine structural heterogeneity, a significant fraction of the apparent multifractal dispersion at finite scales is driven by pre-asymptotic discreteness artifacts — a warning applicable to any discrete lattice study. Second, and equally important, the FSS analysis reveals what single-snapshot analyses cannot: once these transient artifacts are filtered out, the RBIM retains a finite residual dispersion $\Delta H_{\infty}\approx 0.123$ that is absent in the pure model. This residual is the genuine fingerprint of the Griffiths phases. Our results therefore establish systematic FSS as a universal prerequisite for multifractal analysis: it is the only procedure capable of separating these two physically distinct contributions.
@@ VI.4 Generality Beyond the 2D Ising Model
    Systems with quenched disorder (random field Ising, spin glasses, diluted models): Our RBIM results (Fig. 5) demonstrate that even here, finite-size discreteness artifacts heavily inflate the apparent multifractal width. FSS must be strictly applied to determine if any residual broad spectrum genuinely persists asymptotically. Our 2D RBIM analysis demonstrates precisely this capability: the FSS protocol reveals a finite residual dispersion $\Delta H_{\infty}\approx 0.123$ that survives the thermodynamic limit, distinguishing genuine Griffiths complexity from the transient artifacts that dominate finite-size measurements.
@@ VII Conclusion
Our comparative analysis with the Random Bond Ising Model (RBIM) demonstrates both the necessity and the discriminatory power of the proposed FSS protocol. At finite scales, disordered lattices exhibit a broad multifractal dispersion that partially originates from the same pre-asymptotic discreteness artifacts affecting the pure model — a finding that warns against interpreting single-snapshot results as intrinsic physics. However, the FSS analysis reveals a physically meaningful distinction: unlike the pure model, whose spectral width vanishes entirely in the thermodynamic limit, the RBIM retains a finite residual dispersion $\Delta H_{\infty}\approx 0.123$ that persists asymptotically. This residual is the genuine multifractal signature of Griffiths phases, detectable only after the transient artifacts have been filtered out by systematic scaling. Consequently, the FSS protocol is not merely a safeguard against false positives, but an active diagnostic tool: it simultaneously eliminates spurious finite-size broadening and isolates the asymptotic physical signal that single-scale analyses confound with noise..
@@
Beyond theoretical spin models, the methodological vulnerabilities and the FSS protocol established in this work have immediate and critical implications for the analysis of real-world empirical data. MFDFA is extensively applied across diverse physical domains—ranging from the analysis of ocean-atmosphere climate indices (e.g., the Tropical Southern Atlantic Index, TSA) to the study of chaotic rotational dynamics in astrophysics via digitalized light curves (e.g., Saturn’s moon Hyperion).
It is crucial to recognize that empirical time series are fundamentally analogous to the discrete lattices studied here: they are strictly finite in length and subject to digital sampling (discretization). Our results demonstrate that applying MFDFA to a single, finite digitalized record inherently risks generating spuriously broad singularity spectra—often mistakenly interpreted as genuine multifractal complexity—due to the mathematical divergences associated with discrete step-like fluctuations at small scales. Therefore, applying an analog of our proposed scaling protocol (such as systematically analyzing subsets of varying total length N to test for the asymptotic collapse of the H(q) dispersion) is a necessary safeguard. It ensures that the reported multifractality in experimental data reflects true underlying nonlinear multiplicative processes rather than the ubiquitous artifacts of finite digital sampling.